OpenStudy (ebydaqueen19):

An airplane with a speed of 91.9 m/s is climbing upward at an angle of 55.9 ° with respect to the horizontal. When the plane's altitude is 851 m, the pilot releases a package. (a) Calculate the distance along the ground, measured from a point directly beneath the point of release, to where the package hits the earth. (b) Relative to the ground, determine the angle of the velocity vector of the package just before impact.

OpenStudy (raffle_snaffle):

First, draw a picture. =)

OpenStudy (raffle_snaffle):

V = 91.9 m/s Theta = 55.9* h = 851 m Find the horizontal components and the vertical components of the velocity. Then find the time it takes the package to hit the surface. (Hint: time has to be a positive value). Next find the horizontal range. That should get you started.